OpenStudy (idku):

Hi, I am having a problem determining whether or not the statement is true. (Help would be nice.)

OpenStudy (domebotnos):

Hello! what's your statement?

OpenStudy (idku):

\(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\)

OpenStudy (domebotnos):

@jim_thompson5910 you can have this one, buddy.

OpenStudy (idku):

I know how to do \(\color{blue}{\displaystyle(∃x)(∃!y)(x^2-y^2=64)}\). When x=8 (or x=-8), leaving me with only one option, y=0. So that would have been true.

OpenStudy (idku):

The way I read \(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\) is, "there exists a unique (the only) value of y, so that there will be (will exist) an x-value to make the statement true" and I think that this is in fact not true

OpenStudy (idku):

Because, indeed \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64)}\) (for all y, not just for a unique value of y)

OpenStudy (idku):

(I will have to leave for about 30 minutes sorry.)

jimthompson5910 (jim_thompson5910):

I think the commutative law applies in this case https://www.cs.sfu.ca/~ggbaker/zju/math/pred-quant.html scroll to the bottom to the "Nested Quantifiers" section

jimthompson5910 (jim_thompson5910):

Focus on where it says "But it turns out these are equivalent:...i.e. you can swap the same kind of quantifier"

OpenStudy (sshayer):

if x=10,y=6 can you check?

OpenStudy (vuriffy):

100 - 36 = 64. ^

OpenStudy (vuriffy):

I believe the order does count in relevance to the x and y values though.

OpenStudy (sshayer):

i have tried like this . (x+y)(x-y)=64 both can be negative.

OpenStudy (sshayer):

either x=10,y=6 or x=-10,y=-6

OpenStudy (vuriffy):

Yes, which both work in this situation. But, I believe a certain unique value must be found, not completely sure.

OpenStudy (vuriffy):

@idku May you clarify the problem for me a little more? It's been awhile since I touched base on these.

OpenStudy (idku):

@sshayer , so basically you are saying that I can define what my unique valye of y is?

OpenStudy (vuriffy):

As x = 8; x = -8, y must equal 0. In relevance toward, 10 and 6, you are able to configure a problem in which both values in positive and negative terms allow for both formats to succeed in the answer of 64. You may define your unique value of y.

OpenStudy (idku):

In the statement: \(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\) I think you are first to choose the y, and then the x.

OpenStudy (vuriffy):

Yes, in sense of the unique value. Y before the x.

OpenStudy (vuriffy):

I may be incorrect, or in my mind sensing the x may be first, so I will not hold my statement to be completely true.

OpenStudy (idku):

Basically the way I read \(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\) is: "There is only one value of \(\color{blue}{\displaystyle y}\), for which there will exist (at least one value of) \(\color{blue}{\displaystyle x}\) to satisfy \(\color{blue}{\displaystyle x^2-y^2=64}\)."

OpenStudy (idku):

(If "∃x" was first, it would be a lot easier, but I think these aren't equivalent, and I just explained why I think so.)

OpenStudy (vuriffy):

Oh, I see now. Yes, I would believe you must choice the y value first, then the x value after.

OpenStudy (idku):

Really, the following is true: \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64)}\) So, if it is true \(\forall y\), how can I say that there exists a unique value of y (only one value for y) that would satisfy the statement.

OpenStudy (vuriffy):

I apologize, I have forgot much about this, but I can surely inform when y = 0 as you stated as the unique value for y, does allow for at least one value of x to construct toward 64. I wish I could be of further assistance. Maybe @jim_thompson5910 could help you with this part in which you state there exists a unique value of y.

OpenStudy (idku):

Yeah, I am pretty certain the preposition is false. What does \(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\) say? There will exist ONLY one value of \(y\), for which there will exist AT LEAST ONE \(x\) to satisfy \(x^2-y^2=64\). This is not true tho', because FOR ANY VALUE OF \(y\) ((not for a unique one)) , there will exist an \(x\) to satisfy \(x^2-y^2=64\).

OpenStudy (idku):

This is basically the argument I made (just want confirmation, or dispute).

OpenStudy (vuriffy):

Confirmed for one which is not a unique value of y, I would believe so, yes. The preposition would be considered false in my eyes.

OpenStudy (idku):

I will repost my basic argument to save time.

jimthompson5910 (jim_thompson5910):

Ignore my post above. It only works for existential quantifiers (not unique existential quantifiers) You are correct @idku the statement would be false. Simply list 2 or more cases and that will be enough to prove the uniqueness aspect doesn't hold up.

OpenStudy (idku):

The proposition: \(\color{blue}{\displaystyle(∃!y)(∃x)(x^2-y^2=64)}\) My interpretation: There exists a unique \(y\), for which there is going to exist (at least one) \(x\), to satisfy \(x^2-y^2=64\).

OpenStudy (idku):

tnx for the confirmation, Jim.

OpenStudy (idku):

So, in any case, I claimed that in fact \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64)}\) is true, and it is certainly not a unique value of \(y\) for which the rest of the statement holds.

jimthompson5910 (jim_thompson5910):

You don't need to show it for all y You just need more than one y. That's the min amount of work you need to do really.

jimthompson5910 (jim_thompson5910):

But that method will work too

OpenStudy (vuriffy):

y = 0, y = +/- 6, y = +/- 15 are some choices to equal 64 as the result. x = +/- 8, x = +/- 10, x = +/- 17.

OpenStudy (idku):

So, if I actually want to disproof, I just need to specific (numerical) examples?

OpenStudy (idku):

need two**

OpenStudy (vuriffy):

I believe so, as what Jim stated.

OpenStudy (vuriffy):

I gave you 3 cases, just so you could see there was more.

jimthompson5910 (jim_thompson5910):

`So, if I actually want to disproof, I just need two specific (numerical) examples?` correct

OpenStudy (loser66):

To me, if your statement is A, then if \(\neg A\) is False, then A is true. And \(\neg A \) is \(\forall y , \cancel \exists x | x^2-y^2=64\) this statement is False, hence A is true.

OpenStudy (idku):

I lost you on how "for all y" is really a negation to "for a unique y". There are multiple possibilities (also "exists no y such that....)

OpenStudy (loser66):

The original one is "There exists only y", negate it gives you "for all y"

OpenStudy (idku):

And actually, Loser66, I think that \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64)}\) would be true.

OpenStudy (loser66):

If you play guess game, then go with it If you need a proof, I think my way works well. :)

OpenStudy (idku):

For any value of \(y\), there will exist an \(x\) to satisfy \(x^2-y^2=64\). It's an equivalent of \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64) ~~:= ~~}\)\(\color{blue}{\displaystyle(\forall y)(∃x)(|y|=\sqrt{x^2-64})}\)

OpenStudy (idku):

You said that \(\color{blue}{\displaystyle(\forall y)(∃x)(x^2-y^2=64)}\) would not be true, can you show me how then? (And no, I am not guessing, just trying to learn something here:D)

OpenStudy (idku):

Well, he left, I guess I will go with my conclusion of the initial statement being false (and will provide two examples of this).

OpenStudy (idku):

Thanks for participation ;;

OpenStudy (loser66):

My statement is "for all y" "NOT exist x" such that x^2-y^2=64 That is wrong.

jimthompson5910 (jim_thompson5910):

The only time \[\color{blue}{\displaystyle(\exists ! y)(\exists x)(f(x,y)=0)}\] is true is if you have a horizontal line for the graph of f(x,y)=0. However, x^2-y^2 = 64 leads to a hyperbola. So that's another way to look at it.